can you help me solved this question it is 11 11/6 + 14 2/3

Answer by Boreal(15235)

Here is the easiest way yoou can do this.
11 11/6 +14 2/3
Isn't this 11 + 11/6 + 14 +2/3?
Now, add 11 and14=25
Now add 11/6 and 2/3
common denominator 6
11/6 + 4/6 =15/6
This is 2 1/2
you add 2 1/2 to the 25 and get 27 1/2
Alternative
11 11/6= 11 +6/6 +5/6= 12 5/6
Add 12 5/6 to 14 2/3
That is 12 + 14 + 5/6 + 2/3. You can add in any order you wish
26 + 5/6 +2/3 common denominator is 6. You have (5+4)/6= 9/6 = 1 1/2
Add it to 26, and you have 27 1/2
Add the integers, then add the fractions. Don't try to convert these to improper fractions. You would have 77/6 + 44/3. Yes, this may be done, but it is much more difficult.